# Robust Utility-Preserving Text Anonymization Based on Large Language Models

### v1

@@ Appendix A Dataset Settings
To build DB-bio dataset, we sampled data samples from the DBPedia Classes dataset, where each sample consists of the biography, the profile of the described people and the three-level category. We sampled according to the third level category. Specifically, we chose 24 categories, and the number of data samples for each category is shown in fig. 7. Then we manually checked each sample to filter out non-English tokens and examples with a biography longer than 700 words or shorter than 300 words. Finally, we divided the whole dataset into train, validation and test part following the ratio of 8:2:1.
Appendix B Evaluation Metrics
@@
To evaluate our text anonymization method, we focus on two critical aspects: disclosure risk and utility preservation. Disclosure risk is assessed by measuring the success rate (SR) of a strong adversarial LLM in inferring personal information from anonymized text. A lower success rate indicates lower disclosure risk. Different from the P-Evaluator in the anonymization process, a more rigorous case is used in the evaluation set-up, where the ground-truth is mixed with other similar items and the adversarial LLM is prompted to choose one from these items according to the anonymized text. Additionally, we further prompted an LLM to generate the Confidence Scores (CS), evaluating how confidently the anonymized text can be associated with the ground-truth personal information, providing a measure of uncertainty while making inferences
@@
Utility preservation metrics are gauged by the performance of a simple neural network classifier trained on non-anonymized train data but tested on anonymized data, including Accuracy, macro averaged Precision, macro averaged Recall, macro averaged F1 Score, and the classifier’s loss function value indicating classification uncertainty. For the DB-bio dataset, we train a BERT model Devlin et al. (2019) on the train set using validation set for hyper-parameter tuning. In the training process, we set the batch as 16 learning rate as 1e-5. We use the linear learning rate scheduler. We train the model for 20 epochs. For the PersonalReddit dataset, we train a RoBERTa-large Liu et al. (2019) model on the train set and use the test set for hyper-parameter tuning. In the training process, we set the batch as 8, learning rate as 1e-5. We use the linear learning rate scheduler. We train the model for 10 epochs.
Appendix C Implementation Details
C.1 Prompts
@@
Figure 23: The prompt template used to evaluate the confidence score
@@ C.2 Knowledge Distillation
We access GPT-3.5 and GPT-4 through the API provided by Azure. We fine-tuned the two student models using the QLORA method Dettmers et al. (2024). We use the turbo version of GPT-4 for cost saving. For both the SFT and OPT fine-tuning phase, we follow the instruction fine-tuning manner where the instruction "Please anonymize the following biography:" is prepended to the input biography. For the Phi-3 Mini model, we use the released instruction-tuned version of it, we set the learning rate as 2e-4, set the batch size as 4, set the gradient accumulation steps as 4, and the epochs number as 7. The rank and alpha of the QLORA method are set as 32 and 64, respectively. The dropout rate is set as 0.05 For the Llama-3-8b model, we use the released instruction-tuned version of it, we set the learning rate as 1e-4, set the batch size as 4, set the gradient accumulation steps as 4, and the epochs number as 7. The rank and alpha of the QLORA method are set as 32 and 64, respectively. The dropout rate is set as 0.1 For both models, we quantize them with 4 bits. We use the paged adamw 32 bit optimizer and cosine learning rate scheduler. The warmup ratio is set as 0.05. The experiments are conducted on a Nvidia A100 80G GPU.
Appendix D Detailed Related Work
D.1 Text Anonymization
@@
Text anonymization is crucial for protecting privacy in textual data, primarily addressed through natural language processing (NLP) and privacy-preserving data publishing (PPDP) approaches. NLP methods use sequence labeling models trained on manually annotated data to identify and remove pre-defined categories of sensitive information, such as names and phone numbers Hathurusinghe et al. (2021); Francopoulo and Schaub (2020); Adams et al. (2019); Eder et al. (2022); Arranz et al. (2022); Jensen et al. (2021); Kleinberg et al. (2022). NLP approaches typically do not account for non-predefined sensitive information and apply uniform masking to all detected data, lacking flexibility in adjusting the level of anonymization based on disclosure risk.
@@
Privacy-preserving data publishing (PPDP) focuses on developing computational techniques to release data without compromising privacy. The PPDP-based approaches to anonymization is fundamentally privacy-first, enforcing a pre-defined privacy model through various data masking methods such as noise addition or value generalization Chakaravarthy et al. (2008); Cumby and Ghani (2011); Anandan et al. (2012); Sánchez and Batet (2016, 2017). For instance, the well-known k-anonymity privacy model Chakaravarthy et al. (2008) requires that each combination of quasi-identifier attribute values is shared by at least k records in the dataset. However, these methods often impractically assume that sensitive entities are pre-detected or require extensive external data resources to calculate disclosure risk Sánchez and Batet (2016), which limits their practicality in dynamic environments.
@@
The extraordinary capabilities of LLMs significantly influence text anonymization studies. On the one hand, LLMs’ in-context learning ability have diminished the need for manually annotated training data, simplifying domain adaptation in text anonymization tasks Liu et al. (2023); Dou et al. (2023); Albanese et al. (2023). However, the powerful abilities of LLMs also introduce new threats to privacy. Their capacity to semantically infer personal information from texts provided at inference time poses a significant disclosure risk to existing anonymization techniques  Nyffenegger et al. (2024); Staab et al. (2024a); Patsakis and Lykousas (2023), which is largely overlooked both by traditional anonymization methods and emerging LLM-based approaches. In response, a concurrent study by Staab et al. introduced an Adversarial Feedback framework, where one LLM anonymizes texts based on adversarial feedback from another LLM tasked with re-identifying the text, aiming to mitigate re-identification risks from LLMs. Despite its effectiveness in enhancing privacy, this method does not account for the impact on downstream analysis, often compromising the utility of the anonymized text for further use.
D.2 Prompt Optimization with LLMs
@@
To address this limitation, Yang and Li formulated prompt optimization as an evolutionary multi-objective optimization problem. Using an Evolutionary Algorithm, they obtained the Pareto optimal set of prompts, allowing users to choose prompts based on their preferences over multiple criteria. Analogously, the task of text anonymization can also be framed as an multi-objective optimization problem with two conflicting objectives: privacy and utility. Different from prompt optimization, text anonymization explicitly prioritizes privacy and requires a unique optimal anonymization solution for each document. Therefore, we propose to frame text anonymization as a lexicographic optimization problem and leverage LLMs to solve it.

### v2

@@ Appendix D Dataset Settings
In this paper, we assume a threat model where the adversaries utilize an LLM pre-trained or fine-tuned on a corpus containing sensitive information to re-identify personal information from the anonymized free-form text Staab et al. (2024a); Patsakis and Lykousas (2023). To evaluate our method in this threat model efficiently without collecting a sensitive information dataset and training an adversary LLM upon it, we use the following two datasets:
  • DB-bio: Many LLMs are pre-trained on the Wikipedia corpus to get the primary knowledge. Thus, we can assume the celebrity information in Wikipedia as personal information to be anonymized and use existing LLMs that have memorized this information as the adversary LLM to attack anonymization methods. We sampled celebrity biographies from the DBpedia Classes dataset Dan (2019) to build a new dataset DB-bio. Unlike the commonly-used Wiki-bio dataset Lebret et al. (2016) in anonymization studies that lack annotations for downstream tasks, this dataset includes detailed three-level hierarchical category annotations. We use the third-level category labels as occupation classification labels to assess the impact of our anonymization method on this specific downstream task. The name of the person described by the biography is used as the ground-truth personal information.
  • PersonalReddit (PR): Due to the rich existence of the celebrity information in the whole pre-train dataset, off-the-shelf LLMs are proficient at guessing the celebrity information from anonymized text. The evaluation performed on the above dataset can only provide an upper bound of the attack success rate. To further validate the generality of our method, we evaluate it on the PR dataset  Staab et al. (2024a) consisting of 525 human-verified synthetic public Reddit comments and the corresponding user profiles. We use the annotated occupation attribute in the profile as the label of the occupation classification task and anonymize the comments to prevent the identification of other personal attributes like gender and location that are understood by existing LLMs.
General statistics of these two datasets can be seen in table 9.
@@
To build The DB-bio dataset, we sampled data samples from the DBPedia Classes dataset, where each sample consists of the biography, the profile of the described people, and the third-level category. We sampled according to the third level category. Specifically, we chose 24 categories, and the number of data samples for each category is shown in fig. 7. Then we manually checked each sample to filter out non-English tokens and examples with a biography longer than 700 words or shorter than 200 words. Finally, we divided the whole dataset into train, validation, and test parts following the ratio of 8:2:1.
Appendix E Knowledge Distillation Results
Detailed knowledge distillation experiment results can be seen in table 10. Row “Phi-3” and “Llama-3” represent the performance of these models without fine-tuning, and the other rows were results copied from fig. 5 of our paper (We extracted the values used to draw fig. 5 and present them numerically in the table below).
Model SR (s) CS Accuracy Similarity
Phi-3 71.42 74.49 93.22 56.70
-SFT 60.25 65.73 97.48 62.12
-DPO 54.15 55.90 92.15 53.90
Llama-3 69.38 71.56 95.40 58.92
-SFT 56.90 59.97 96.33 60.29
-DPO 51.03 53.78 93.28 54.02
Table 10: Knowledge distillation experiment results.
Appendix F Evaluation Metrics
@@
To evaluate our text anonymization method, we focus on two critical aspects: disclosure risk and utility preservation. Disclosure risk is assessed by measuring the success rate (SR) of a strong adversarial LLM in inferring personal information from anonymized text. A lower success rate indicates lower disclosure risk. Different from the P-Evaluator in the anonymization process, a more rigorous case is used in the evaluation set-up, where the ground truth is mixed with other similar items and the adversarial LLM is prompted to choose one from these items according to the anonymized text. Additionally, we further prompted an LLM to generate the Confidence Scores (CS), evaluating how confidently the anonymized text can be associated with the ground-truth personal information, providing a measure of uncertainty while making inferences.
@@
Utility preservation metrics are gauged by the performance of a simple neural network classifier trained on non-anonymized train data but tested on anonymized data, including Accuracy, macro averaged Precision, macro averaged Recall, macro averaged F1 Score, and the classifier’s loss function value indicating classification uncertainty. For the DB-bio dataset, we train a BERT model Devlin et al. (2019) on the train set using the validation set for hyper-parameter tuning. In the training process, we set the batch as 16 learning rate as 1e-5. We use the linear learning rate scheduler. We train the model for 20 epochs. For the PersonalReddit dataset, we train a RoBERTa-large Liu et al. (2019) model on the train set and use the test set for hyperparameter tuning. In the training process, we set the batch as 8 and, the learning rate as 1e-5. We use the linear learning rate scheduler. We train the model for 10 epochs.
Due to the difference between the original test set and the anonymized test set, there exists an out-of-distribution (OOD) problem that will affect the performance of our evaluation classifier trained on the original training set. The general viewpoint on OOD issues is that when the mismatch between the training and test datasets is less significant, the neural networks should perform better than otherwise. This is consistent with what we demonstrated in the utility preservation experiments. If the anonymization method keeps the original text as much as possible, then the OOD issue is expected to be less significant , and the performance drop of the neural network classifier on the anonymized test dataset should be less too. As shown in Table 1, named entity recognition (NER)-based methods replaced too many entities and thus got the highest performance drop. While RUPTA achieved the lowest performance drop due to its preservation of the original text. Similarly, as shown in Table 2, NER-based methods suffered from recognizing implicit personally identifiable information, they almost didn’t anonymize anything, thus they got the lowest performance drop. Since the goal of non-training metrics is also to compare the original and the anonymized text to evaluate whether the anonymization methods keep the original text as much as possible, we expect the trends on the non-training metrics will be similar to what have observed in the paper. Therefore, we didn’t use other non-training metrics to compare the original and anonymized text directly.
Appendix G Implementation Details
G.1 Main experiments
For AF and our method, we set the maximum iteration number, namely the value of $T$, as 5 to make a fair comparison. We set the value of $K$ as 10 to obtain the main results of the experiment. In the experiment, we use the Microsoft Azure platform to access GPT-3.5 and GPT-4, the API version is 2023-05-15. The version of GPT-3.5 is 0301. The version of GPT-4 is 0613.
G.2 Prompts
@@
Figure 23: The prompt template used to evaluate the confidence score.
@@ G.3 Knowledge Distillation
We access GPT-3.5 and GPT-4 through the API provided by Azure. We fine-tuned the two student models using the QLORA method Dettmers et al. (2024). We use the turbo version of GPT-4 for cost savings. For both the SFT and OPT fine-tuning phases, we follow the instruction fine-tuning manner where the instruction "Please anonymize the following biography:" is prepended to the input biography. For the Phi-3 Mini model, we use the released instruction-tuned version of it, we set the learning rate as 2e-4, set the batch size as 4, set the gradient accumulation steps as 4, and the epochs number as 7. The rank and alpha of the QLORA method are set as 32 and 64, respectively. The dropout rate is set as 0.05. For the Llama-3-8b model, we use the released instruction-tuned version of it, we set the learning rate as 1e-4, set the batch size as 4, set the gradient accumulation steps as 4, and the epochs number as 7. The rank and alpha of the QLORA method are set as 32 and 64, respectively. The dropout rate is set as 0.1. For both models, we quantize them with 4 bits. We use the paged adamw 32 bit optimizer and cosine learning rate scheduler. The warmup ratio is set as 0.05. The experiments are conducted on a Nvidia A100 80G GPU.
Appendix H Detailed Related Work
H.1 Text Anonymization
@@
Text anonymization is crucial for protecting privacy in textual data, primarily addressed through natural language processing (NLP) and privacy-preservation data publishing (PPDP) approaches. NLP methods use sequence labeling models trained on manually annotated data to identify and remove pre-defined categories of sensitive information, such as names and phone numbers Hathurusinghe et al. (2021); Francopoulo and Schaub (2020); Adams et al. (2019); Eder et al. (2022); Arranz et al. (2022); Jensen et al. (2021); Kleinberg et al. (2022). NLP approaches typically do not account for non-predefined sensitive information and apply uniform masking to all detected data, lacking flexibility in adjusting the level of anonymization based on disclosure risk.
@@
Privacy-preserving data publishing (PPDP) focuses on developing computational techniques to release data without compromising privacy. The PPDP-based approaches to anonymization are fundamentally privacy-first, enforcing a pre-defined privacy model through various data masking methods such as noise addition or value generalization Chakaravarthy et al. (2008); Cumby and Ghani (2011); Anandan et al. (2012); Sánchez and Batet (2016, 2017). For instance, the well-known k-anonymity privacy model Chakaravarthy et al. (2008) requires that each combination of quasi-identifier attribute values is shared by at least k records in the dataset. However, these methods often impractically assume that sensitive entities are pre-detected or require extensive external data resources to calculate disclosure risk Sánchez and Batet (2016), which limits their practicality in dynamic environments.
@@
The extraordinary capabilities of LLMs significantly influence text anonymization studies. On the one hand, LLMs’ in-context learning ability has diminished the need for manually annotated training data, simplifying domain adaptation in text anonymization tasks Liu et al. (2023); Dou et al. (2023); Albanese et al. (2023). However, the powerful abilities of LLMs also introduce new threats to privacy. Their capacity to semantically infer personal information from texts provided at inference time poses a significant disclosure risk to existing anonymization techniques  Nyffenegger et al. (2024); Staab et al. (2024a); Patsakis and Lykousas (2023), which is largely overlooked both by traditional anonymization methods and emerging LLM-based approaches. In response, a concurrent study by Staab et al. introduced an Adversarial Feedback framework, where one LLM anonymizes texts based on adversarial feedback from another LLM tasked with re-identifying the text, aiming to mitigate re-identification risks from LLMs. Despite its effectiveness in enhancing privacy, this method does not account for the impact on downstream analysis, often compromising the utility of the anonymized text for further use.
H.2 Prompt Optimization with LLMs
@@
To address this limitation, Yang and Li formulated prompt optimization as an evolutionary multi-objective optimization problem. Using an Evolutionary Algorithm, they obtained the Pareto optimal set of prompts, allowing users to choose prompts based on their preferences over multiple criteria. Analogously, the task of text anonymization can also be framed as a multi-objective optimization problem with two conflicting objectives: privacy and utility. Different from prompt optimization, text anonymization explicitly prioritizes privacy and requires a unique optimal anonymization solution for each document. Therefore, we propose to frame text anonymization as a lexicographic optimization problem and leverage LLMs to solve it.